OpenStudy (anonymous):

What is the 22nd term of the arithmetic sequence where a1 = 8 and a9 = 56?

OpenStudy (anonymous):

134 142 150 158

OpenStudy (anonymous):

@campbell_st @ranga

OpenStudy (anonymous):

answers a22= 134 a41=-145

OpenStudy (campbell_st):

well the formula is \[a_(n) = a_{1} + (n - 1)\times d\] you'll need this to find the common difference d \[56 = 8 + (9 -1) \times d\] solve for d then use the formula again with n = 22 to find the 22nd term

OpenStudy (anonymous):

so 56 = 16(d) Divide both sides to get d? @campbell_st

OpenStudy (campbell_st):

well you have 56 = 8 + 8d sbutract 8 from both sides 48 = 8d now solve for d

OpenStudy (anonymous):

d=6

OpenStudy (campbell_st):

thats right so you now need \[a_{22} = 8 + (22 -1) \times 6\]

OpenStudy (anonymous):

so then its a(n)=a1+(n−1)×d a22=8+(22−1)×6 a22 = 8 + 126 a22 = 134

OpenStudy (campbell_st):

well done

OpenStudy (anonymous):

Thank you! @campbell_st Seriously, I didn't understand the whole equation thing until I got your help on this one. I seriously just had that click!

OpenStudy (campbell_st):

glad to help